OpenStudy (anonymous):

6xy^2/7y^3 / 15x2y/5x^2

OpenStudy (anonymous):

Is that 15x^2y or 15x*2y?

OpenStudy (anonymous):

15x^2y

OpenStudy (anonymous):

\[\frac{6xy^2}{7y^3} \div \frac{15x^2y}{5x^2}\]\[=\frac{6xy^2}{7y^3} \times \frac{5x^2}{15x^2y}\]\[=\frac{30x^{1+2}y^2}{105x^2y^{3+1}} \]\[=\frac{2x^{3}y^2}{7x^2y^{4}} \]